OpenStudy (anonymous):

please Help : Which Ratio Forms A propotion with 5 over 6 6 over 5 , 20 over 24, 20 over 30, or 24 over 20

OpenStudy (anonymous):

5/6 = 5/6 times 1 = 5/6 times 4/4 = 20/24

OpenStudy (gw2011):

If you cross multiply 5/6 with 20/24, you will get (20)(6)=120 and (5)(24)=120. Therefore, 20/24 is the only one in which the proportion comes out to 120=120

OpenStudy (anonymous):

thank you

OpenStudy (gw2011):

You're welcome